patrick is 19 years older than his son pat. in 10 years patrick will be twice as old as pat. how old are they?
Found 2 solutions by ikleyn, addingup:
Answer by ikleyn(52798)

.
Patrick is 19 years older than his son pat. in 10 years Patrick will be twice as old as pat. how old are they?
~~~~~~~~~~~~~~~~~~~~~~

In 10 years they will be 19 and 38. (which is obvious/evident).

Hence, Patrick is 38-10 = 28 years now.


Answer by addingup(3677)

Let's call the son x and the father y
y = x+19 we'll use this value for y next
y+10 = 2(x+10)
x+19+10 = 2x+20
x+29 = 2x+20 subtract x and 20 on both sides
9 = x
or
x = 9 this is how old Pat is.
>>>>>>>>>>>>>>>>>>>>>>>>
Check:
the father is 19 years older than the son:
9+19 = 28
In 10 years he will be twice as old as the son:
28+10 = 38 is the father. And the son:
2*(9+10) = 38 Correct
:
John